[Brahms-l] Shift report 20050502 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon May 02 2005 - 07:05:30 EDT
Shifter leader was SB
Shifters were 


1. Beam was not aborted by C-AD  inspite of the Fire Alarm at BRAHMS.
   Data taking was not inturrepted and we had 3.25 beam hours. 
2. FS@8 deg @2442B, MRS@45 deg @1050A
   trig2(11.7k),trig3(64k),trig4(97k),
   trig5(92k),trig5(28k)
3. No detector problems, all detectors in PARK mode
   All magnets in STANDBY

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon May 2 07:05:56 2005

This archive was generated by hypermail 2.1.8 : Mon May 02 2005 - 07:06:02 EDT